# LEPTODERMIC CORRECTIONS TO THE TOV EQUATIONS AND NUCLEAR ASTROPHYSICS WITHIN THE EFFECTIVE SURFACE APPROXIMATION

### v1

@@ Abstract
The macroscopic model for a neutron star (NS) as a liquid drop at the equilibrium is used to extend the Tolman-Oppenheimer-Volkoff (TOV) equations taking into account the gradient terms responsible for the system surface. The parameters of the Schwarzschild metric in the spherical case are found with these surface corrections in the leading (zero) order of the leptodermic approximation $a/R<<1$, where $a$ is the NS effective-surface (ES) thickness, and $R$ is the effective NS radius. The energy density $\mathcal{E}$ is considered in a general form including the functions of the particle number density and of its gradient terms. The macroscopic gravitational potential $\Phi(\rho)$ is taken into account in the simplest form as expansion in powers of $\rho-\overline{\rho}$, where $\overline{\rho}$ is the saturation density, up to second order, in terms of its contributions to the separation particle energy and incompressibility. Density distributions $\rho$ across the NS ES in the normal direction to the ES, which are derived in the simple analytical form at the same leading approximation, was used for the derivation of the modified TOV (MTOV) equations by accounting for their NS surface corrections. The MTOV equations are analytically solved at first order and the results are compared with the standard TOV approach of the zero order.
@@ I INTRODUCTION
The macroscopic effective-surface approach (ESA) wilets ; strtyap ; tyapin ; strmagbr ; strmagden ; MS09 ; BM13 ; BM15 taking into account the gradient terms of the energy density was applied MS24 for the neutron stars. Following the works of Tolman, Oppenheimer, and Volkoff (TOV) RT39 ; OV39 we consider the neutron star (NS) as a liquid drop at equilibrium under the gravitational, and other realistic forces; see also the Tolman and Rowlinson books RT87 ; RW82 . As shown in Ref. RT87 , the TOV equations were derived OV39 from the equations of the General Relativistic Theory (GRT) assuming the spherical symmetry and the macroscopic approximation to the energy-momentum tensor. The expression for the energy-momentum tensor was essentially used in the TOV derivations for a macroscopic system as the perfect liquid drop LLv2 ; LLv6 . Many applications in terms of the Schwarzschild metric solutions in the spherical case can be found in Refs. LLv2 ; WF88 ; CB97 ; SH06 ; Ko08 ; FCPG13 ; AB14 ; BC15 ; LH19 ; SBL23 . For the basic relations of the nuclear and neutron liquid-drop models to NS properties we should also mention a pioneer work BBP71 done by Baym, Bethe, and Pethick. They discussed the liquid matter drop with the leptodermic property having a sharp decrease of the particle number density in a relatively small edge region considered as its surface. See also Refs. LLv2 ; RT87 ; OV39 ; SBL23 ; LP01 ; HP01PRL ; HP01PRC ; PH00Poland ; ABCG09 ; PFCPS13 ; FCPG13 ; FG23 ; DFG23 ; HPY07 ; LH19 ; ST04 for the description of the finite NS dense matter systems in terms of the leptodermic approximation, $a/R\ll 1$, where $a$ is the crust thickness, and $R$ is the mean curvature radius of the ES. Clear specific definitions and complete updated results for the energy density with density gradient (surface) terms and for equations of the infinite matter state with many inter-particle forces in the non-relativistic and relativistic cases can be found in the recent review Ref. SBL23 .
@@
Figure 1: Particle number density, $\rho$, in units of the saturation value, $\overline{\rho}$, as function of the radial variable $r$ (in km) for a neutron star in a simple compressed liquid-drop model at the stable equilibrium. Solid line is related to the asymmetric solution, Eq. (32) [the asterisk means the equation taken from Table 1 of Ref. MS24 , line “1”], which is the leading approximation over a small leptodermic parameter, $a/R\ll 1$. Dashed line presents the same but for the Wilets symmetric solution wilets ; MS24 (line “2”). Parameters, for example: the effective radius $R=10$ km, diffuseness of the NS crust $a=1$ km, and temperature $T=0$; see, e.g., Refs. ST04 ; HPY07 ; MS24 . The dots ES1 and ES2 show the ES at the effective radius $R$ for the solid and dashed lines, respectively (from Ref. MS24 ).
@@
(8) $\Phi$ Macroscopic gravitational potential
@@
Table 1: Equations of Ref. MS24 (the first column), marked by the asterisk in the text, are presented for the quantity with symbols in the second column, and explained in the third column. The vdW-Skyrme is the van der Waals and Skyrme inter-particle interaction.
@@
In this work we derive the modified TOV (MTOV) equations by taking into account the gradient surface corrections within the leptodermic approximation Refs. LJ23 ; LP01 ; HP01PRL ; HP01PRC ; PH00Poland ; ABCG09 ; PFCPS13 ; FCPG13 ; FG23 ; DFG23 ; HPY07 ; LH19 ; MS24 . We consider first a dense one-component system as in Ref. MS24 , and take the energy density functional $\mathcal{E}$ in a rather general form, linear in $(\nabla\rho)^{2}$ terms, with the coefficients which are smooth functions of the particle density $\rho$. In such a system, the particle density $\rho$ in the leading order approximation is function of the radial coordinate with exponentially decreasing behavior from almost a constant saturation value inside of the dense system to that of zero through the NS effective surface in a relatively small crust range $a$; see, e.g., Fig. 1. The ES is defined as the inflection points of spatial coordinates, i.e., with density gradients maximum. To obtain the analytical solutions for the particle number density and Equation of State (EoS), we use the leptodermic approximation. In this approximation, the gravitational forces can be taken into account through the macroscopic gravitational potential $\Phi$ determined by the second order expansion over deflection of the density $\rho$ from its saturation value $\overline{\rho}$ (see, e.g., Ref. MS24 ).
@@
Within the leptodermic approximation, $a/R\ll 1$, simple and accurate solutions of many nuclear and liquid-drop problems involving the particle number density distributions were obtained by using the ES approach for nuclei wilets ; strtyap ; tyapin ; strmagbr ; strmagden ; MS09 ; BM13 ; BM15 ; see also Ref. MS24 for applications to the neutron stars. The ESA exploits the property of saturation of the nuclear particle density $\rho$ inside of the system, which is a characteristic feature of dense systems such as molecular systems, liquid drops (amorphous solids), nuclei, and presumably, NSs. The realistic energy-density distribution per particle is minimal at a certain saturation density $\overline{\rho}$, corresponding approximately to the infinite dense matter bete . As a result, relatively narrow edge region exists in finite nuclei or NS (crust) in which the density drops sharply from its almost central value to zero. We assume here that the part inside of the system far from the ES can be changed a little (saturation property of the dense system as hydrostatic liquid drop, nucleus, or NS in the final evolution state).
@@
The coordinate system related to the effective surface is defined in such a way that one of the spatial coordinates ($\xi$) is the distance from a given point to the ES, and others ($\eta$) are tangent to the ES. This nonlinear coordinate system $\xi,\eta$ is conveniently used in a region near the nuclear and NS edge. They allow for an easy extraction of the relatively large terms in density distribution equations for the condition of a variational equilibrium. This condition means that the variation of the total energy $E$ over the density $\rho$ is zero under the constraints which fix some integrals of motion beyond the energy $E$ by the Lagrange method. The Lagrange multipliers are determined by these constraints within the local energy-density theory, in particular, the extended Thomas-Fermi (ETF) approach from nuclear, dense molecular, and metallic-cluster physics brguehak ; brbhad . Neglecting the other smaller-order perpendicular- and all parallel-to-ES contributions, sum of such terms leads to a simple one-dimensional equation (in special local coordinates with the coordinate normal-to-surface $\xi$). Such an equation mainly determines approximately the density distribution across the diffused surface layer. The latter is of the relatively small order over the ratio of the diffuseness parameter $a$ to the (mean curvature) ES radius $R$, $a/R<<1$, for sufficiently heavy systems. A small parameter, $a/R$, of the expansion within the ES approach can be used for analytical solving the variational problem for the minimum of the energy density functional per volume with constraints for a fixed particle number, and other integrals of motion, such as angular momentum, quadrupole deformation, etc. When this edge distribution of the density is known, the leading static and dynamic density distributions which correspond to the diffused surface conditions can be easily constructed. To do so, one has to determine the dynamics of the effective surface which is coupled to the volume dynamics of the density by certain liquid-drop model boundary conditions bormot ; strmagbr ; magstr . Our ESA is based on the catastrophe theory for solving differential equations with a small parameter of the order of $a/R$ as the coefficient in front of the high order derivatives in normal to the ES direction Ti52 . A relatively large change of the density $\rho$ on a small distance $a$ with respect to the curvature radius $R$ takes place for a finite liquid-matter drop (nuclei, molecular water drops, neutron stars). We emphasize that the macroscopic ESA theory MS24 allows to consider the deformed and superdeformed systems for which the same coordinate system $\xi,\eta$ can be used locally near the ES. Inside of such dense systems, the density $\rho$ is changed a little around constant saturation density $\overline{\rho}$. Therefore, including also the second order expansion of the gravitational potential near $\overline{\rho}$, one obtains essential effects of the surface capillary pressure.
@@
The accuracy of the ESA was checked in Ref. strmagden for the case of static and dynamic nuclear physics by comparing the results with the existing nuclear theories like Hartree-Fock (HF) vauthbrink and ETF brguehak ; brbhad , based on the Skyrme forces vauthbrink ; skyrme ; barjac ; ringshuk ; blaizot ; brguehak ; gramvoros ; krivin ; CB97 ; CB98 , but for the simplest case without spin-orbit and asymmetry terms of the energy density functional. The direct variational principle for finding numerically the parameters of the tested particle number density functions in simple forms of the Woods-Saxon-like in Ref. BB03 or their powers (Ref. kolsan ) were applied by using the realistic Skyrme energy functional CB97 ; CB98 ; see also recently published variational ETF approach KS18 ; KS20 . The extension of the ES approach to the nuclear isotopic symmetry and spin-orbit interaction has been done in Refs. MS09 ; BM13 ; BM15 . The Swiatecki derivative terms of the symmetry energy for heavy nuclei myswann69 ; myswiat85 ; danielewicz2 ; vinas1 ; vinas2 ; vinas4 ; vinas5 ; Pi09 were taken into account within the ESA in Ref. BM15 . The discussions of the progress in nuclear physics and astrophysics within the relativistic local density approach, can be found in reviews Refs. SBL23 ; NVR11 ; see also Refs. CH08 ; Pi21 ; Pe23 .
@@
In the present work we apply the effective surface approach for the neutron stars of Ref. MS24 , based on Refs. strmagden ; MS09 ; BM13 ; BM15 of nuclear physics, to derive the MTOV equations taking into account the surface leptodermic corrections. In Sect. II we present the TOV derivations at the zero order approach. In Sect. III we derive the MTOV equations with accounting for the first order surface corrections to the TOV equations. The derivations of the analytical solutions to the MTOV equations are obtained too. Sect. IV is devoted to the discussions of the results. These results are summarized in the conclusion section V. Some details of the calculations are given in Appendix A.
II Tolman-Oppengeimer-Volkoff approach
@@
where $c$ is the speed of light. Functions $\lambda=\lambda(r)$ and $\nu=\nu(r)$ are given by the differential equations (see Ref. RT87 at zero cosmological constant $\Lambda$). First, outside of the system on big distances from masses, or for relatively small gravitational forces RT87 ; LLv2 , one can derive these equations using the macroscopic properties of a dense spherical system for the energy-momentum tensor. Transforming then the radial coordinate $r$ of the spherical four-coordinate system $r,\theta,\phi$, and $t$, due to the GRT invariance, for the line element in this space, Eq. (1), one obtains the simple differential equations for $\lambda$ and $\nu$ of the Schwarzschild metric for the region inside of the liquid-drop system (see, Refs. RT87 ; LLv2 ),
$\frac{8\pi G}{c^{4}}P=e^{-\lambda}\left(\frac{\nu^{\prime}}{r}+\frac{1}{r^{2}}\right)-\frac{1}{r^{2}}~{},$ (2)
@@
$\frac{dP}{dr}=-\frac{1}{2}\left(\mathcal{E}+P\right)\nu^{\prime}~{},$
@@
Using the substitution $e^{-\lambda}=f$, one can solve the second equation in Eq. (2), as a linear first-order differential equation over $f(r)$, in terms of the quadratures. Then, the derivative $\nu^{\prime}$ can be found from the first equation of Eq. (2). Integrating it over $r$, one obtains the Schwarzschild metric (1) in terms of these solutions for $\lambda(r)$ and $\nu(r)$. Substituting the derivative $\nu^{\prime}$ into the third equation of Eq. (2), one finally obtains the famous TOV equations RT39 ; OV39 ; LJ23 ,
@@
The integrations are terminated when $P=0$, which defines the surface $r=R$. A specified value of the central pressure $P_{0}=P(r=0)$ determines the NS volume mass $M=m(r=R)$. Notice that the surface corrections to the NS volume mass were obtained in Ref. MS24 .
As shown in Ref. RT87 , taking approximately into account the step-function density, $\rho=\overline{\rho}$ for $r<R$ inside of the system, and $\rho=0$ for $r>R$ outside of it, one can solve TOV equations (II) explicitly analytically (see Refs. RT87 ; MS24 ). (For instance, $\overline{\rho}$ is the constant density $\rho_{00}$ using units $G=c=1$, inside of the system in the notations of Tolman). Our leading order solutions for the density $\rho$ are obtained in Ref. MS24 by using the ESA within the leading leptodermic approximation; see Eqs. (31) for more general inter-particle interaction and (32) for the specific vdW-Skyrme forces. They have a sharp transition from the saturation to zero value, which agrees with above mentioned approximation at zero order of the leptodermic expansion in small parameter $a/R\ll 1$ (see Fig. 1). Taking also into account the first boundary condition in Eq. (4), the second equation in Eq. (II) can be integrated analytically,
@@
$\mathcal{E}_{0}=\mathcal{E}(\rho=\overline{\rho})~{}\Theta\left(\overline{\rho}-\rho\right)=\mathcal{A}(\overline{\rho})\Theta(R-r),$ (6)
and $\overline{\rho}$ is the saturation approximation for the density $\rho$. We introduced here the Heaviside step function, $\Theta(x)$, $\Theta(x)=1$ for $x\geq 0$ and 00 for $x<0$. The function $\mathcal{A}(\rho)$ in Eq. (6) is given by Eq. (9) of Ref. MS24 for $\mathcal{A}(\rho)$ at the saturation density $\rho=\overline{\rho}$; see Table 1. The density $\overline{\rho}$ and energy density $\mathcal{E}_{0}$ can be taken approximately as those found in Ref. MS24 ; see Eqs. (3), (9), (10), and (11), including the macroscopic gravitational field. Formally, one can consider the expression (9) for $\mathcal{E}_{0}(\overline{\rho})$ inside of the system [or Eq. (16)] as the macroscopic EoS additional to the TOV equations (II). Therefore, the total mass $M$ neglecting the surface effects and the influence of the Schwarzschild metric curvature, is given within this zero order ESA approximation by
@@
where $R_{\rm S}$, $A_{\rm S}$, and $B_{\rm S}$ are the specific integration constants, and $r<R_{\rm S}$. This metric, Eq. (II), is nonsingular at $r=0$, in contrast to the original Schwarzschild metric RT87 ; LLv2 , which is valid outside of the gravitating system. Therefore, it can be applied for the part of the space inside of this system (see Ref. RT87 ). A smooth match of the external and internal Schwarzschild metrics through the effective radius, $r=R$, is assumed to be carried out. The radius $R_{\rm S}$ in Eq. (II) is expressed in terms of the energy density $\mathcal{E}$ inside of the system, $\mathcal{E}_{0}$, far away from the ES as
@@
Figure 2: Contour plots for the pressure $P(r)$, Eq. (11), in units of the central value, $P(r=0)$ as function of the radial coordinate $r$ and the dimensionless parameter $A_{\rm S}$ of the Schwarzschild metric, Eq. (II). The numbers in squares show the values of this pressure. White color presents regions where we have indetermination, infinity by infinity, with the finite limit 1 at $r\rightarrow 0$. Red color shows negative values of the ratio $P(r)/P(r=0)$ [a positive pressure $P(r)$]. The value “0.00” displays the zero value in the horizontal and vertical coordinate lines on right of plots. The effective NS radius $R=10$ km is the same as in Fig.  1 (from Ref. MS24 ).

### v2

@@ Abstract
The macroscopic model for a neutron star (NS) as a liquid drop at the equilibrium is used to extend the Tolman-Oppenheimer-Volkoff (TOV) equations taking into account the gradient terms responsible for the system surface. The parameters of the Schwarzschild metric in the spherical case are found with these surface corrections to the known leading (zero) order of the leptodermic approximation $a/R<<1$, where $a$ is the NS effective-surface (ES) thickness, and $R$ is the effective NS radius. The energy density $\mathcal{E}$ is considered in a general form including the functions of the particle number density and of its gradient terms. The macroscopic gravitational component $\Phi(\rho)$ of the energy density is taken into account in the simplest form as expansion in powers of $\rho-\overline{\rho}$, where $\overline{\rho}$ is the saturation density, up to second order, in terms of its contributions to the separation particle energy and incompressibility. Density distributions $\rho$ across the NS ES in the normal direction to the ES, which are derived in the simple analytical form at the same leading approximation, was used for the derivation of the modified TOV (MTOV) equations by accounting for their NS surface corrections. The MTOV equations are analytically solved at first order and the results are compared with the standard TOV approach of the zero order.
@@ I INTRODUCTION
The macroscopic effective-surface approach (ESA) wilets ; strtyap ; tyapin ; strmagbr ; strmagden ; MS09 ; BM13 ; BM15 taking into account the gradient terms of the energy density was applied MM24 for the neutron stars. Following the works of Tolman, Oppenheimer, and Volkoff (TOV) RT39 ; OV39 we consider the neutron star (NS) as a finite, dense and perfect liquid drop at equilibrium under the gravitational, and other realistic forces; see also the Tolman book RT87 (chapt. 7, sect. 96). As shown in Ref. RT87 , the TOV equations were derived OV39 from the equations of the General Relativistic Theory (GRT) assuming the spherical symmetry and the macroscopic approximation for the energy-momentum tensor. The expression for the energy-momentum tensor was essentially used in the TOV derivations for a macroscopic system as the perfect liquid drop (see Refs. LLv2 (chapts. 11 and 12), LLv6 (chapt. 1, sects. 1-3; chapt. 7, sect. 61; chapt. 15, sect. 133) and Rowlinson-Widow RW82 (chapt. 1). Many applications of the TOV result in terms of the Schwarzschild metric solutions, Ref. KS16 , for spherically symmetric NSs can be found in Refs. LLv2 ; WF88 ; CB97 ; ST04 ; SH06 ; HPY07 ; Ko08 ; PFCPS13 ; AB14 ; BC15 ; LH19 ; SBL23 (sect. 9 in Ref. ST04 and sect. 3 in Ref. HPY07 ). For the basic relations of the nuclear and neutron liquid-drop models to NS properties we should also mention a pioneer work BBP71 done by Baym, Bethe, and Pethick. They discussed the liquid matter drop with the leptodermic property having a sharp decrease of the particle number density $\rho$ in a relatively small edge region considered as its surface. Notice, in the TOV derivations, the authors assumed that the outer and inner Schwarzschild metrics have to be matched at a definite NS radius that looks meaningful for a leptodermic behavior of the density $\rho$. For a good hint to the nuclear astrophysics; see also, e.g., Refs. LP01 ; HP01PRL ; HP01PRC ; PH00Poland ; CH08 ; ABCG09 ; FCPG13 ; LH19 ; FG23 ; DFG23 ; LJ23 ; Pe23 ; Pe24 ; XV24 . Clear specific definitions and complete updated results for the energy density with density gradient (surface) terms and for equations of the infinite matter state with many inter-particle forces in the non-relativistic and relativistic cases can be found in the recent review Ref. SBL23 .
@@
Figure 1: Particle number density, $\rho$, in units of the saturation value, $\overline{\rho}$, as function of the radial variable $r$ (in km) for a neutron star in a simple compressed liquid-drop model at the stable equilibrium. Solid line is related to the asymmetric solution, Eq. (32) [the asterisk means the equation taken from Table 1 of Ref. MM24 , line “1”], which is the leading approximation over a small leptodermic parameter, $a/R\ll 1$. Dashed line presents the same but for the Wilets symmetric solution wilets ; MM24 (line “2”). Parameters, for example: the effective radius $R=10$ km, diffuseness of the NS crust $a=1$ km, and temperature $T=0$; see, e.g., Refs. ST04 ; HPY07 ; MM24 . The dots ES1 and ES2 show the ES at the effective radius $R$ for the solid and dashed lines, respectively (from Ref. MM24 ).
@@
(8) $\Phi$ Macroscopic gravitational field
@@
Table 1: Equations of Ref. MM24 (the first column), marked by the asterisk in the text, are presented for the quantity with symbols in the second column, and explained in the third column. The vdW-Skyrme is the van der Waals and Skyrme inter-particle interaction.
Notice that the Schwarzschild radius $R_{S}$ RT39 ; OV39 ; RT87 can be considered as a constant which determines the upper limit of the NS radius and can be used for the transition to the black holes if the particle density $\rho$ has a saturation property $\rho\approx\overline{\rho}$ inside of the NS. Therefore, the energy density $\mathcal{E}$ which is the argument of $R_{S}$, $R_{S}\propto(\mathcal{E}(\overline{\rho}))^{-1/2}$, has also the corresponding saturation property. Otherwise, we have to modify the TOV equations for the pressure $P(r)$ to agree with the equation of state (EoS) which would allow a smooth behavior of the density $\rho(r)$. As the TOV equation OV39 is used along with the EoS in a lot of astrophysical articles on the NS, it would be logical to agree the arguments for the specific derivations of both these equations. This is the main motivation of our and TOV macroscopic approaches. Notice also that of course, in nuclear astrophysics we have to take into account the strong gravitational forces as in the derivations of the TOV equations, in contrast to the nuclear physics. Our suggestion for these coordinations does not exclude more microscopic EoSs. However, it requires the corresponding modification of the TOV equations which, unfortunately, makes the problem to be more complicate, far away from the present analytical approach.
@@
In this work we derive the modified TOV (MTOV) equations by taking into account the gradient surface components of the energy density which are suggested basically, e.g., in Refs. BBP71 ; LP01 ; HP01PRL ; HP01PRC ; PH00Poland ; HPY07 ; CH08 ; ABCG09 ; PFCPS13 ; FCPG13 ; LH19 ; FG23 ; DFG23 ; SBL23 ; LJ23 ; Pe23 ; Pe24 ; XV24 ; MM24 . Within the leptodermic approximation MM24 based on such gradient density components of the energy density, one can consider first analytically a dense one-component system. In these calculations we take the energy density functional $\mathcal{E}(\rho)$ in a rather general form, as linear in $(\nabla\rho)^{2}$ terms, with the coefficients which are smooth functions of the particle density $\rho$. In such a system, the particle density $\rho$ in the leading order approximation is function of the radial coordinate with exponentially decreasing behavior from almost a constant saturation value inside of the dense system to that of zero through the NS effective surface in a relatively small crust range $a$; see, e.g., Fig. 1. The ES is defined as the inflection point of spatial coordinates, i.e., with density gradients maximum. In order to obtain the analytical solutions for the particle number density and EoS, we use the leptodermic macroscopic approximation. In this approximation, the gravitational forces can be taken into account through the macroscopic gravitational field $\Phi$ determined by the second order expansion over deflection of the density $\rho$ from its saturation value $\overline{\rho}$ (see Ref. MM24 ).
@@
Within the leptodermic approximation, $a/R\ll 1$, simple and accurate solutions of many nuclear and liquid-drop problems involving the particle number density distributions were obtained by using the ES approach for nuclei wilets ; strtyap ; tyapin ; strmagbr ; strmagden ; MS09 ; BM13 ; BM15 ; see also Ref. MM24 for applications to the neutron stars. The ESA exploits the property of saturation of the nuclear particle density $\rho$ inside of the system, which is a characteristic feature of dense systems such as molecular systems RW82 ; vdW , liquid drops (amorphous solids), nuclei, and presumably, NSs, in agreement with the TOV derivations. The realistic energy-density distribution per particle is minimal at a certain saturation density $\overline{\rho}$, corresponding approximately to the infinite dense matter bete . As a result, relatively narrow edge region exists in finite nuclei or NS (crust) in which the density drops sharply from its almost central value to zero. We assume here that the part inside of the system far from the ES can be changed a little (saturation property of the dense system as hydrostatic liquid drop, nucleus, or presumably, NS in a final evolution state).
@@
The coordinate system related to the effective surface is defined in such a way that one of the spatial coordinates ($\xi$) is the distance from a given point to the ES, and others ($\eta$) are tangent to the ES. This nonlinear coordinate system $\xi,\eta$ is conveniently used in a region near the nuclear and NS edge. They allow for an easy extraction of the relatively large terms in density distribution equations for the condition of a variational equilibrium. This condition means that the variation of the total energy $E$ over the density $\rho$ is zero under the constraints which fix some integrals of motion beyond the energy $E$ by the Lagrange method. The Lagrange multipliers are determined by these constraints within the local energy-density theory, in particular, the extended Thomas-Fermi (ETF) approach from nuclear, dense molecular, and metallic-cluster physics brguehak ; brbhad . Neglecting the other smaller-order perpendicular- and all parallel-to-ES contributions, sum of such terms leads to a simple one-dimensional equation (in special local coordinates with the coordinate normal-to-surface $\xi$). Such an equation mainly determines approximately the density distribution across the diffused surface layer. The latter is of the relatively small order over the ratio of the diffuseness parameter $a$ to the (mean curvature) ES radius $R$, $a/R<<1$, for sufficiently heavy systems. A small parameter, $a/R$, of the expansion within the ES approach can be used for analytical solving the variational problem for the minimum of the energy density functional per volume with constraints for a fixed particle number, and other integrals of motion, such as angular momentum, quadrupole deformation, etc. When this edge distribution of the density is known, the leading static and dynamic density distributions which correspond to the diffused surface conditions can be easily constructed. To do so, one has to determine the dynamics of the effective surface which is coupled to the volume dynamics of the density by certain liquid-drop model boundary conditions bormot ; strmagbr ; magstr . Our ESA is based on the catastrophe theory for solving differential equations with a small parameter of the order of $a/R$ as the coefficient in front of the high order derivatives in normal to the ES direction Ti52 . A relatively large change of the density $\rho$ on a small distance $a$ with respect to the curvature radius $R$ takes place for a finite liquid-matter drop (nuclei, molecular water drops, neutron stars). We emphasize that the macroscopic ESA theory MM24 allows to consider the deformed and super-deformed systems for which the same coordinate system $\xi,\eta$ can be used locally near the ES. Inside of such dense systems, the density $\rho$ is changed a little around constant saturation density $\overline{\rho}$. Therefore, including also the second order expansion of the gravitational field near $\overline{\rho}$, one obtains essential effects of the surface capillary pressure which were originally studied for molecular systems by van der Waals; see Refs. RW82 ; vdW .
@@
The accuracy of the ESA was checked in Ref. strmagden for the case of static and dynamic nuclear physics by comparing the results with the existing nuclear theories like Hartree-Fock (HF) vauthbrink and ETF brguehak ; brbhad , based on the Skyrme forces vauthbrink ; skyrme ; barjac ; ringshuk ; blaizot ; brguehak ; gramvoros ; krivin ; CB97 ; CB98 , but for the simplest case without spin-orbit and asymmetry terms of the energy density functional. The direct variational principle for finding numerically the parameters of the tested particle number density functions in simple forms of the Woods-Saxon-like in Ref. brbhad or their powers (Ref. kolsan ) were applied by using the realistic Skyrme energy functional CB97 ; CB98 ; see also recently published variational ETF approach KS18 ; KS20 . The extension of the ES approach to the nuclear isotopic symmetry and spin-orbit interaction has been done in Refs. MS09 ; BM13 ; BM15 . The Swiatecki derivative terms of the symmetry energy for heavy nuclei myswann69 ; myswiat85 ; danielewicz2 ; vinas1 ; vinas2 ; vinas4 ; vinas5 ; Pi09 were taken into account within the ESA in Ref. BM15 . The discussions of the progress in nuclear physics and astrophysics within the relativistic local density approach, can be found in review articles, Refs. SBL23 ; NVR11 ; see also Refs. CH08 ; Pi21 ; Pe23 ; Pe24 .
@@
In the present work we apply the effective surface approach for the neutron stars of Ref. MM24 , based on Refs. strmagden ; MS09 ; BM13 ; BM15 of nuclear physics, to derive the MTOV equations taking into account the surface leptodermic corrections. In Sect. II we present the TOV derivations at the zero order approach. In Sect. III we derive the MTOV equations with accounting for the first order surface corrections to the TOV equations. The derivations of the analytical solutions to the MTOV equations are obtained too. Section IV is devoted to the discussions of the results. These results are summarized in the conclusion section V. Some details of the calculations are given in Appendix A.
II Tolman-Oppenheimer-Volkoff approach
@@
where $c$ is the speed of light. Functions $\lambda=\lambda(r)$ and $\nu=\nu(r)$ are given by the differential equations (see Ref. RT87 at zero cosmological constant $\Lambda$). First, outside of the system on big distances from masses, or for relatively small gravitational forces RT87 ; LLv2 , one can derive these equations using the macroscopic properties of a dense spherical system for the energy-momentum tensor. Transforming then the radial coordinate $r$ of the spherical four-coordinate system $r,\theta,\phi$, and $t$, due to the GRT in-variance, for the line element in this space, Eq. (1), one obtains the simple differential equations for $\lambda$ and $\nu$ of the Schwarzschild metric for the region inside of the liquid-drop system (see, Refs. RT87 ; LLv2 ),
$\frac{8\pi G}{c^{4}}P=e^{-\lambda}\left(\frac{\nu^{\prime}}{r}+\frac{1}{r^{2}}\right)-\frac{1}{r^{2}}~,$
@@
$\frac{dP}{dr}=-\frac{1}{2}\left(\mathcal{E}+P\right)\nu^{\prime}~,$ (2)
@@
Using the substitution $e^{-\lambda}=f$, one can solve the second equation in Eq. (II), as a linear first-order differential equation over $f(r)$, in terms of the quadratures. Then, the derivative $\nu^{\prime}$ can be found from the first equation of Eq. (II). Integrating it over $r$, one obtains the Schwarzschild metric (1) in terms of these solutions for $\lambda(r)$ and $\nu(r)$. Substituting the derivative $\nu^{\prime}$ into the third equation of Eq. (II), one finally obtains the famous TOV equations RT39 ; OV39 ; LJ23 ,
@@
The integrations are terminated when $P=0$, which defines the surface $r=R$. A specified value of the central pressure $P_{0}=P(r=0)$ determines the NS volume mass $M=m(r=R)$. Notice that the surface corrections to the NS volume mass were obtained in Ref. MM24 .
As shown in Ref. RT87 , taking approximately into account the step-function density, $\rho=\overline{\rho}$ for $r<R$ inside of the system, and $\rho=0$ for $r>R$ outside of it, one can solve TOV equations (II) explicitly analytically (see Refs. RT87 ; MM24 ). (For instance, $\overline{\rho}$ is the constant density $\rho_{00}$ using units $G=c=1$, inside of the system in the notations of Tolman). Our leading order solutions for the density $\rho$ are obtained in Ref. MM24 by using the ESA within the leading leptodermic approximation; see Eqs. (31) for more general inter-particle interaction and (32) for the specific vdW-Skyrme forces. They have a sharp transition from the saturation to zero value, which agrees with above mentioned approximation at zero order of the leptodermic expansion in small parameter $a/R\ll 1$ (see Fig. 1). Taking also into account the first boundary condition in Eq. (4), the second equation in Eq. (II) can be integrated analytically,
@@
$\mathcal{E}_{0}=\mathcal{E}(\rho=\overline{\rho})~\Theta\left(\rho-\overline{\rho}\right)=\mathcal{A}(\overline{\rho})\Theta(R-r),$ (6)
and $\overline{\rho}$ is the saturation approximation for the density $\rho$. We introduced here the Heaviside step function, $\Theta(x)$, $\Theta(x)=1$ for $x\geq 0$ and 0 for $x<0$. The function $\mathcal{A}(\rho)$ in Eq. (6) is given by Eq. (9) of Ref. MM24 for $\mathcal{A}(\rho)$ at the saturation density $\rho=\overline{\rho}$; see Table 1. The density $\overline{\rho}$ and energy density $\mathcal{E}_{0}$ can be taken approximately as those found in Ref. MM24 ; see Eqs. (3), (9), (10), and (11), including the macroscopic gravitational field. Formally, one can consider the expression (9) for $\mathcal{E}_{0}$ inside of the system [or Eq. (16)] as the macroscopic EoS additional to the TOV equations (II). Therefore, the total mass $M$ neglecting the surface effects and the influence of the Schwarzschild metric curvature, is given within this zero order ESA approximation by
@@
where $R_{\rm S}$, $A_{\rm S}$, and $B_{\rm S}$ are the specific integration constants, and $r<R_{\rm S}$. This metric, Eq. (II), is non-singular at $r=0$, in contrast to the original Schwarzschild metric RT87 ; LLv2 , which is valid outside of the gravitating system. Therefore, it can be applied for the part of the space inside of this system (see Ref. RT87 ). A smooth joining of the external and internal Schwarzschild metrics through the effective radius, $r=R$, is assumed to be carried out. The Schwarzschild radius $R_{\rm S}$ in Eq. (II) is expressed in terms of the energy density $\mathcal{E}$ inside of the system, $\mathcal{E}_{0}$, far away from the ES as
@@
Figure 2: Contour plots for the pressure $P(r)$, Eq. (11), in units of the central value, $P(r=0)$ as function of the radial coordinate $r$ and the dimensionless parameter $A_{\rm S}$ of the Schwarzschild metric, Eq. (II). The numbers in squares show the values of this pressure. White color presents regions where we have in-determination, infinity by infinity, with the finite limit 1 at $r\rightarrow 0$. Red color shows negative values of the ratio $P(r)/P(r=0)$ [a positive pressure $P(r)$]. The value “0.00” displays the zero value in the horizontal and vertical coordinate lines on right of plots. The effective NS radius $R=10$ km is the same as in Fig.  1 (from Ref. MM24 ).